OpenStudy (anonymous):

How do you write a polynomial function with rational coefficients in standard form with the given zeros?

OpenStudy (anonymous):

and the zeros are: -3, 0, 0, 5

OpenStudy (mertsj):

f(x)=(x+3)(x)(x)(x-5)=x^2(x^2-2x-15)=x^4-2x^3-15x^2

OpenStudy (anonymous):

Thanks =)

OpenStudy (mertsj):

yw